The prelim2e package^{∗}

Martin Schröder

Crüsemannallee 3 D-28213 Bremen MS@Dream.HB.North.DE (INTERNET)

1997/05/12

Abstract

This package allows the marking of (preliminary) versions of a document on the output.

Contents

1 Introduction 1

2 Usage 1

3 Options 2 4 Required packages 2

5 Putting more things at the bottom 2

1 Introduction

This package allows the marking of (preliminary) versions of a document. This is done using the command \PrelimText, whose expansion is added below the footer of every page of a document (look at the bottom of this page for an example).

2 Usage

Simply using this package via \usepackage{prelim2e} produces a text in the form of “Preliminary version – June 9, 1997”.

\PrelimText The text is produced by the commands \PrelimText and \PrelimWords, which

\PrelimWords can be changed via \renewcommand* or by using options at the \usepackage

command (see section 3).

The footer for this documentation was produced by changing \PrelimWords to

     \renewcommand*{\PrelimWords}{%
        \textsf{prelim2e} package --
        Version \fileversion{} --
        Documentation \LaTeX{}ed%
        }

1

∗The version number of this file is v1.20, last revised 1997/05/12.

prelim2e package – Version v1.20 – Documentation L^{A}TEXed – June 9, 1997 – 13:54

3 Options

The package has the following options:

draft If this option is used a text appears below the normal pagebody. It is the default.

final If this option is used prelim2e produces no text.

english This sets the text to “Preliminary version”. It is the default.

french This sets the text to “Version préliminaire”. It does not use the french or

New feature 1997/05/12

babel package.

german This sets the text to “Vorläufige Version”. It does not use the german or babel package.

New description time This adds the time at the beginning of the current L^{A}TEX to the text. The 1997/05/12 time is produced using the \thistime command, which may be provided by

the scrtime package. If you don’t use the scrtime package (e. g. by specifying the scrtime option), you must provide it yourself or by some other package.

scrtime This loads the scrtime package (part of the KOMA-Script bundle), which

New feature 1997/05/12

provides the \thistime command used by the time option. The scrtime option implies the time option.

All other options are passed to the scrtime package if the scrtime option is selected.

4 Required packages

The package requires the following packages:

everyshi It is used to implement the setting of the text below the normal pagebody.

New description scrtime It is used to typeset the current time and is needed if the scrtime option 1997/05/12 is selected.

5 Putting more things at the bottom

Another package you might like to use with prelim2e is vrsion. This allows the definition and maintenance of a version number like 3.14159 within L^{A}TEX, which can be put at the bottom of every page using prelime2e.

prelim2e package – Version v1.20 – Documentation L^{A}TEXed – June 9, 1997 – 13:54